Solve without using determinant form A.B×C= (Axi+Ayj+Azk). (Bxi+Byj+Bzk)× (Cxi+Cyj+Czk)


1
Expert's answer
2021-08-17T17:31:36-0400

"B\times C=\begin{bmatrix}
 \hat{i}&\hat{j}&\hat{k} \\B_x&B_y&B_z \\
 C_x&C_y&C_z 
\end{bmatrix}"


"B\times C=\hat{i}(B_yC_z-B_zC_y)-\hat{j}(B_xC_z-B_zC_x)+\hat{k}(B_xC_y-B_yC_x)"

"A.(B\times C)=(A_x\hat{i}+\hat{j}A_y+\hat{k}A_z).(\hat{i}(B_yC_z-B_zC_y)-\hat{j}(B_xC_z-B_zC_x)+\hat{k}(B_xC_y-B_yC_x))"

"A.(B\times C)=A_x(B_yC_z-B_zC_y)-A_y(B_xC_z-B_zC_x)+A_z(B_xC_y-B_yC_x)"